A rectangular garden has a perimeter of 34 yards. Fencing across the length costs $2 per yard and fencing along the width costs $1 per yard. The total cost of the fencing is $58. Find the length and width of the rectangle.
Answer by checkley77(12844)

2L+2W=34
2L=34-2W
2*2L+1*2W=58 REPLACE 2L IN THIS EQUATION WITH (34-2W) & SOLVE FOR THE WIDTH.
2(34-2W)+2W=58
68-4W+2W=58
-2W=58-68
-2W=-10
W=-10/-2
W=5 ANS. FOR THE WIDTH.
2L=34-2*5
2L=34-10
2L=24
L=24/2
L=12 ANS. FOR THE LENGTH.
PROOFS:
2*12+2*5=34
24+10=34
34=34
2*2L+1*2W=58
4*12+2*5=58
48+10=58
58=58